# Moment maps and stability of holomorphic submersions

### v1

@@ 4. Stability and optimal symplectic connections
The following result [37, Lemma 3.7] describes the subspace of $V_{\pi}$ which parametrises relatively K-semistable fibrations with smooth fibres admitting a degeneration to a relatively cscK fibration.
@@ 4.1. A finite-dimensional moment map on fibrations
For $\xi\in\mathfrak{k}_{\pi,v}$, let $\sigma_{\xi}$ be the corresponding fibrewise holomorphic vector field and $h_{\xi}$ be its fibrewise holomorphy potential with respect to the relative Kähler metric $(\omega,I)$. If we perturb $\omega$ to $\omega+i\partial\bar{\partial}\varphi$, for some Kähler potential $\varphi\in\mathcal{K}_{E}(I)$, then [44, Lemma 4,10] the function
@@
In this expression, $E(\varphi,x_{y})$ is the vector bundle of fibre holomorphy potentials with respect to the Kähler structure $(\omega+i\partial\bar{\partial}\varphi,\Phi(x_{y}))$ and $\mathcal{K}_{E}^{2,p+2}(g_{y})$ is the $W^{2,p+2}(g_{y})$-completion of $\mathcal{K}_{E}(\Phi(x_{y}))$. As recalled in §3.2, the differential of $\mathcal{F}_{y}$ at $(\xi,0)$ with respect to $\varphi$ is [39, §5.3]
@@
We next prove that $L_{y}$ has a bounded right inverse. As a differential operator on the vector bundle $E(x_{y})$, the term $-\mathcal{A}_{y}^{*}\mathcal{A}_{y}(\psi)-h_{\xi}-\sigma_{\xi}(\psi)$ is linear [39, §5.3], so $L_{y}$ has a bounded right inverse if and only if $\mathcal{R}_{y}^{*}\mathcal{R}_{y}$ does. As in the proof of Theorem 2.10, the bound follows from a Poincaré type inequality and Schauder estimates. Recall that by definition
@@
and analogously for the holomorphic operators $\partial_{0}$, ¯$\bar{\partial}_{0}$ with respect to the complex structure $\Phi(0)$. These splittings define two Laplace operators on $X$ [34, §2]:
@@
The map $\vartheta$ defined as
$\langle\vartheta(y),v\rangle=\int_{\mathcal{U}_{y}}\langle\tau,v\rangle|_{y}\left(p_{E(x_{y})}\left(\Theta(\omega_{y},\Phi(x_{y}))\right)+\frac{1}{2}\nu_{x_{y}}\right)\omega_{y}^{m}\wedge\omega_{B}^{n}$ (4.4)
@@
A point $y\in V_{\pi}^{+}$ is a zero of the moment map $\vartheta$ if and only if
@@
The upshot is that we can re-define $\omega_{\mathcal{U}}$ to take into account the change in the Kähler potential $\varphi$ and that the zeroes of the moment map $\vartheta$ obtained from $\omega_{\mathcal{U}}$ are optimal symplectic connections.
@@ 4.2. Optimal symplectic connections on stable deformations
Let $Y_{w}\to B$ be a deformation of $Y\to B$, with $w\in V_{\pi}^{+}$. Assume that $Y\to B$ admits an optimal symplectic connection and that $Y_{w}\to B$ is polystable. Then there is a $w_{1}\in V_{\pi}^{+}$ in the $K_{\pi,v}^{\mathbb{C}}$-orbit of $w$ such that $\vartheta(w_{1})=0$.
@@
Assume that $w$ has discrete stabiliser. From Proposition 2.5, the moment map flow of $\vartheta$ starting at $w$ converges to a point $w_{\infty}$, which belongs to $V_{\pi}$ by Proposition 2.4. Assume that $w_{\infty}\notin K_{\pi,v}^{\mathbb{C}}\cdot w$. By Proposition 2.6 we can find $\xi\neq 0$ in $\mathfrak{k}_{\pi,v}$ and $\widetilde{w}\in V_{\pi}$ such that
@@
If $w$ has non-discrete stabiliser, we project orthogonally to a maximal torus in the stabiliser as explained in §2.1. Let $\mathfrak{k}_{T^{\perp}}$ be the Lie algebra (2.2) defined from $\mathfrak{k}_{\pi,v}$, and $K_{T^{\perp}}$ the corresponding group. Arguing as before, we obtain $\widetilde{w}\in K_{T^{\perp}}\cdot w$ such that $\vartheta_{T^{\perp}}(\widetilde{w})=0$. From [44, Theorem 5.25], it follows that there exists $w_{1}\in K_{T^{\perp}}^{\mathbb{C}}\cdot\widetilde{w}$ such that $w_{1}$ is a critical point for $\lVert\theta\rVert^{2}$. This implies [44, Lemma 5.23] that $\theta(w_{1})\in\left(\mathfrak{k}_{\pi,v}\right)_{w_{1}}$. Since $\widetilde{w}$ is in the same orbit as $w$, so is $w_{1}$, so the function $\theta(w_{1})$ also belongs to the Lie algebra $\left(\mathfrak{k}_{\pi,v}\right)_{w}$ of the stabiliser of $w$. Therefore stability implies that
@@ Appendix A Stability of deformations of cscK manifolds
The holomorphic vector field induced by $\xi$ defines a test configuration $(\mathcal{Y},\mathcal{L}^{\prime})$ for $(Y,L)$, represented on $V$ by the closure of the orbit of $\widetilde{x}$ for the action of the one-parameter subgroup of $K^{\mathbb{C}}$ induced by $\xi$. From Theorem 2.12,
$\langle\mu(\widetilde{x}),\xi\rangle=\int_{U_{\tilde{x}}}(\mathrm{Scal}(\omega_{\tilde{x}},\Phi(\widetilde{x}))-\widehat{S})f\omega_{\tilde{x}}^{n},$
where $f$ is the holomorphy potential on $X$ induced by $\xi$. In particular $\langle\mu(\widetilde{x}),\xi\rangle$ is equal to the Futaki invariant of the vector field $\sigma_{\xi}$ induced by $\xi$. By Theorem 2.8 this is, up to a constant, equal to the opposite of the Donaldson-Futaki invariant of the test configuration $(\mathcal{Y},\mathcal{L}^{\prime})$. K-polystability of $Y$ then implies that
@@
If the stabiliser of $x$ is not discrete, we can project the moment map orthogonally to the stabiliser, as explained in §2.1. Then arguing as before, we find a zero $x_{\infty}$ of $\mu_{T^{\perp}}$. From [44, Theorem 5.25], it follows that there exists $x_{1}\in K_{T^{\perp}}^{\mathbb{C}}\cdot x_{\infty}$ such that $x_{1}$ is a critical point for $\lVert\mu\rVert^{2}$. This implies [44, Lemma 5.23] that $\mu(x_{1})\in\mathfrak{k}_{x_{1}}$. Since $(\mathcal{U}_{x_{\infty}},\mathcal{L}_{x_{\infty}})$ is isomorphic to $(Y,L^{\prime})$, so is $(\mathcal{U}_{x_{1}},\mathcal{L}_{x_{1}})$, so the function $\mu(x_{1})$ also belongs to the Lie algebra $\mathfrak{k}_{x}$ of the stabiliser of $x$. Therefore stability implies that
@@
Our proof of Theorem A.1 fixes a mistake in Székelyidhi’s original proof [43, Proposition 8]. In fact, Székelyidhi’s argument relies on a lower bound on the norm of the differential of $\mu$ along the infinitesimal vector field that is uniform in a ball around the origin; this uniform bound cannot exist, because the differential of $\mu$ vanishes at the fixed points, which form a linear subspace. So [43, Proposition 9] cannot be applied. A similar issue is present also in [47, Proposition 3.3.2], where the argument would need a uniform bound on the Hessian of the norm-squared of $\nu$ which cannot hold for the same reason, and in [38, Theorem 1.27]. A related mistake can be found in [2, §2.4], where the techniques require the Kähler property but the finite-dimensional reduction gives a moment map with respect to a symplectic form that is not Kähler.
While our approach applies generally to perturbation problems, a workaround to this issue was developed in certain cases by Inoue in the case of Kähler-Ricci solitons [27, §4.1.2] and by Fan in the case of Higgs bundles [18] using global techniques specific to those problems.
@@
The GIT polystability of $v$ implies that, by the Hilbert–Mumford criterion [22, Theorem 12.2] the weight of the action on $\widetilde{H}^{1}$ of the one-parameter subgroup of $K$ defined by $\xi$

### v2

@@ 4. Stability and optimal symplectic connections
The following result [39, Lemma 3.7] describes the subspace of $V_{\pi}$ which parametrises relatively K-semistable fibrations with smooth fibres admitting a degeneration to a relatively cscK fibration.
@@ 4.1. A finite-dimensional moment map on fibrations
For $\xi\in\mathfrak{k}_{\pi,v}$, let $\sigma_{\xi}$ be the corresponding fibrewise holomorphic vector field and $h_{\xi}$ be its fibrewise holomorphy potential with respect to the relative Kähler metric $(\omega,I)$. If we perturb $\omega$ to $\omega+i\partial\bar{\partial}\varphi$, for some Kähler potential $\varphi\in\mathcal{K}_{E}(I)$, then [47, Lemma 4,10] the function
@@
In this expression, $E(\varphi,x_{y})$ is the vector bundle of fibre holomorphy potentials with respect to the Kähler structure $(\omega+i\partial\bar{\partial}\varphi,\Phi(x_{y}))$ and $\mathcal{K}_{E}^{2,p+2}(g_{y})$ is the $W^{2,p+2}(g_{y})$-completion of $\mathcal{K}_{E}(\Phi(x_{y}))$. As recalled in §3.2, the differential of $\mathcal{F}_{y}$ at $(\xi,0)$ with respect to $\varphi$ is [41, §5.3]
@@
We next prove that $L_{y}$ has a bounded right inverse. As a differential operator on the vector bundle $E(x_{y})$, the term $-\mathcal{A}_{y}^{*}\mathcal{A}_{y}(\psi)-h_{\xi}-\sigma_{\xi}(\psi)$ is linear [41, §5.3], so $L_{y}$ has a bounded right inverse if and only if $\mathcal{R}_{y}^{*}\mathcal{R}_{y}$ does. As in the proof of Theorem 2.10, the bound follows from a Poincaré type inequality and Schauder estimates. Recall that by definition
@@
and analogously for the holomorphic operators $\partial_{0}$, ¯$\bar{\partial}_{0}$ with respect to the complex structure $\Phi(0)$. These splittings define two Laplace operators on $X$ [36, §2]:
@@
The map $\theta$ defined as
$\langle\theta(y),v\rangle=\int_{\mathcal{U}_{y}}\langle\tau,v\rangle|_{y}\left(p_{E(x_{y})}\left(\Theta(\omega_{y},\Phi(x_{y}))\right)+\frac{1}{2}\nu_{x_{y}}\right)\omega_{y}^{m}\wedge\omega_{B}^{n}$ (4.4)
@@
A point $y\in V_{\pi}^{+}$ is a zero of the moment map $\theta$ if and only if
@@
The upshot is that we can re-define $\omega_{\mathcal{U}}$ to take into account the change in the Kähler potential $\varphi$ and that the zeroes of the moment map $\theta$ obtained from $\omega_{\mathcal{U}}$ are optimal symplectic connections.
@@ 4.2. Optimal symplectic connections on stable deformations
Let $Y_{w}\to B$ be a deformation of $Y\to B$, with $w\in V_{\pi}^{+}$. Assume that $Y\to B$ admits an optimal symplectic connection and that $Y_{w}\to B$ is polystable. Then there is a $w_{1}\in V_{\pi}^{+}$ in the $K_{\pi,v}^{\mathbb{C}}$-orbit of $w$ such that $\theta(w_{1})=0$.
@@
Assume that $w$ has discrete stabiliser. From Proposition 2.5, the moment map flow of $\theta$ starting at $w$ converges to a point $w_{\infty}$, which belongs to $V_{\pi}$ by Proposition 2.4. Assume that $w_{\infty}\notin K_{\pi,v}^{\mathbb{C}}\cdot w$. By Proposition 2.6 we can find $\xi\neq 0$ in $\mathfrak{k}_{\pi,v}$ and $\widetilde{w}\in V_{\pi}$ such that
@@
If $w$ has non-discrete stabiliser, we project orthogonally to a maximal torus in the stabiliser as explained in §2.1. Let $\mathfrak{k}_{T^{\perp}}$ be the Lie algebra (2.2) defined from $\mathfrak{k}_{\pi,v}$, and $K_{T^{\perp}}$ the corresponding group. Arguing as before, we obtain $\widetilde{w}\in K_{T^{\perp}}\cdot w$ such that $\theta_{T^{\perp}}(\widetilde{w})=0$. From [47, Theorem 5.25], it follows that there exists $w_{1}\in K_{T^{\perp}}^{\mathbb{C}}\cdot\widetilde{w}$ such that $w_{1}$ is a critical point for $\lVert\theta\rVert^{2}$. This implies [47, Lemma 5.23] that $\theta(w_{1})\in\left(\mathfrak{k}_{\pi,v}\right)_{w_{1}}$. Since $\widetilde{w}$ is in the same orbit as $w$, so is $w_{1}$, so the function $\theta(w_{1})$ also belongs to the Lie algebra $\left(\mathfrak{k}_{\pi,v}\right)_{w}$ of the stabiliser of $w$. Therefore stability implies that
@@ Appendix A Stability of deformations of cscK manifolds
The $\Phi(\widetilde{x})$-holomorphic vector field induced by $\xi$ defines a test configuration $(\mathcal{Y},\mathcal{L}^{\prime})$ for $(Y,L^{\prime})$, represented on $V$ by the closure of the orbit of $x$ for the action of the one-parameter subgroup of $K^{\mathbb{C}}$ induced by $\xi$. From Theorem 2.12,
$\langle\mu(\widetilde{x}),\xi\rangle=\int_{\mathcal{U}_{\tilde{x}}}(\mathrm{Scal}(\omega_{\tilde{x}},\Phi(\widetilde{x}))-\widehat{S})f\omega_{\tilde{x}}^{n},$
where $f$ is the holomorphy potential with respect to $\Phi(\widetilde{x})$ induced by $\xi$. In particular $\langle\mu(\widetilde{x}),\xi\rangle$ is equal to the Futaki invariant of the vector field $\sigma_{\xi}$ induced by $\xi$. By Theorem 2.8 this is, up to a constant, equal to the opposite of the Donaldson-Futaki invariant of the test configuration $(\mathcal{Y},\mathcal{L}^{\prime})$. K-polystability of $Y$ then implies that
@@
If the stabiliser of $x$ is not discrete, we can project the moment map orthogonally to the stabiliser, as explained in §2.1. Then arguing as before, we find a zero $x_{\infty}$ of $\mu_{T^{\perp}}$. From [47, Theorem 5.25], it follows that there exists $x_{1}\in K_{T^{\perp}}^{\mathbb{C}}\cdot x_{\infty}$ such that $x_{1}$ is a critical point for $\lVert\mu\rVert^{2}$. This implies [47, Lemma 5.23] that $\mu(x_{1})\in\mathfrak{k}_{x_{1}}$. Since $(\mathcal{U}_{x_{\infty}},\mathcal{L}_{x_{\infty}})$ is isomorphic to $(Y,L^{\prime})$, so is $(\mathcal{U}_{x_{1}},\mathcal{L}_{x_{1}})$, so the function $\mu(x_{1})$ also belongs to the Lie algebra $\mathfrak{k}_{x}$ of the stabiliser of $x$. Therefore stability implies that
@@
Our proof of Theorem A.1 fixes a mistake in Székelyidhi’s original proof [46, Proposition 8]. In fact, Székelyidhi’s argument relies on a lower bound on the norm of the differential of $\mu$ along the infinitesimal vector field that is uniform in a ball around the origin; this uniform bound cannot exist, because the differential of $\mu$ vanishes at the fixed points, which form a linear subspace. So [46, Proposition 9] cannot be applied. A similar issue is present also in [50, Proposition 3.3.2], where the argument would need a uniform bound on the Hessian of the norm-squared of $\nu$ which cannot hold for the same reason, and in [40, Theorem 1.27]. A related mistake can be found in [3, §2.4], where the techniques require the Kähler property but the finite-dimensional reduction gives a moment map with respect to a symplectic form that is not Kähler.
While our approach applies generally to perturbation problems, a workaround to this issue was developed in certain cases by Inoue in the case of Kähler-Ricci solitons [29, §4.1.2] and by Fan in the case of Higgs bundles [20] using global techniques specific to those problems.
@@
The GIT polystability of $v$ implies that, by the Hilbert–Mumford criterion [24, Theorem 12.2] the weight of the action on $\widetilde{H}^{1}$ of the one-parameter subgroup of $K$ defined by $\xi$